OpenStudy (anonymous):

The area of room A is (5x^2 – 7x – 6) ft^2. Room B has an area of (8x^2 + 6x – 1) ft^2. If room A is the larger room, how much greater is the area of room A than the area of room B?

OpenStudy (linyu):

\[5x ^{2}-7x-6>8x ^{2}+6x-1\] \[-3x ^{2}-13x-5>0\] \[3x ^{2}+13x+5<0\] \[\Delta=13^{2}-4\times3\times5=109\] \[x=\frac{ -13-\sqrt{109} }{ 2\times3 }=-3.9\] \[x=\frac{ -13+\sqrt{109} }{ 2\times3 }=-0.4\] the difference of area is\[-3x ^{2}-13x-5\] \[-3(-3.9)^{2}-13(-3.9)-5=-45.63+50.7-5=0.07\] \[-3(-0.4) ^{2}-13(-0.4)-5=-0.48+5.2-5=-0.28\] since room A is larger than B so 0.07 is the answer. Very apologize if I'm wrong.

OpenStudy (linyu):

3rd

OpenStudy (linyu):

sure

OpenStudy (linyu):

I'm not quite understand the question but \[5^{1+10}\]\[5^{2+9}\]\[5^{3+8}\]